# The Fermi surface of RuO2 measured by quantum oscillations

### v1

@@ Abstract
The metallic oxide RuO2 had emerged as a promising altermagnet candidate, owing to reports of this material hosting antiferromagnetic ordering accompanied by a spin-split electronic band structure characteristic of time-reversal symmetry-breaking. However, recent studies have robustly questioned this orthodoxy. In particular, surface-sensitive probes have detected unusual surface states that may exhibit several characteristics expected for an altermagnet, challenging the determination of the intrinsic magnetic and electronic properties of the bulk. Here we map the Fermi surface of pristine single-crystalline RuO2. By measuring magnetic quantum oscillations of a bulk thermodynamic property, our study resolves the electronic structure present in the bulk of RuO2. Several Fermi sheets are discerned, with heavy effective quasiparticle masses up to five times that of the bare electron mass, indicative of strong electronic correlations. We compare our measurements with the predictions for altermagnetic and nonmagnetic Fermi surfaces deduced from density functional theory calculations. The quantum oscillatory frequency spectra correspond very poorly to the profile expected for the case of altermagnetism; by contrast, they correspond well to the nonmagnetic scenario. Our findings place strong constraints on the bulk magnetic properties of RuO2.
@@ I Introduction
The recently proposed phenomenon of an altermagnetic (AM) phase of matter – characterized by a compensated collinear magnetic structure with a spin polarization that periodically alternates through both the crystal structure in real space and the electronic band structure in reciprocal space – has attracted widespread attention [1, 2, 3, 4, 5, 6, 7, 8, 9, 10]. In addition to being of great fundamental interest – constituting a novel, distinct class of magnetic order – there are numerous properties of AM materials that make them highly desirable for technological applications [11, 12]. For example, their time-reversal symmetry-broken electronic band structures may, as per ferromagnetic materials, manifest spin-polarized currents accompanied by the anomalous Hall effect [4, 3]. Furthermore, attempts to miniaturize ferromagnetic components within integrated circuits often suffer deleterious effects from stray fields causing interference between neighboring elements. By contrast, the lack of a net magnetization M for AM components could enable them to be spatially positioned much closer together in next-generation spintronic circuitry.
@@ II Experimental techniques
For torque magnetometry measurements, samples were aligned by Laue diffractometry and affixed to a BeCu cantilever using a low-temperature adhesive varnish. This technique constitutes a bulk-sensitive thermodynamic measurement that provides a highly sensitive probe of the anisotropy in a material’s magnetic susceptibility tensor. Due to its high fidelity, torque magnetometry is often deployed for dHvA effect measurements in high magnetic fields [30, 31, 32, 33]. In HFML-EMFL the magnetic torque $\bm{\tau}=\mu_{0}(\textbf{M}\crossproduct\textbf{H})$ was measured by a General Radio analogue capacitance bridge, and calibrated to absolute units using an Andeen-Hagerling (AH) digital capacitance bridge. At SECUF, we measured $\tau$ directly with an AH bridge. In both sets of torque experiments, samples were rotated in-situ, with the orientation calibrated by a Hall sensor. For contactless resistivity measurements, samples were also aligned by Laue diffractometry, and affixed to hand-wound copper coils by low-temperature adhesive varnish. TDO and PDO measurements were obtained utilizing a similar methodology to that outlined in ref. [34]. At WHMFC, samples were also rotated in-situ, with the orientation determined by a pick-up coil.
Figure 1: Angle-dependent quantum oscillation measurements. Background-subtracted magnetic torque ($\Delta\tau$) measured at incremental angles in the [001]—[100] rotation plane recorded at (a) SECUF and (b) HFML-EMFL. (c) Quantum oscillations in the contactless resistivity measured at WNHMFC, also in the [001]—[100] plane. (d) Fast Fourier transforms (FFTs) of the data from panel a. These spectra have been renormalized and offset for ease of presentation. (e) Angular distribution of quantum oscillatory frequency spectra, determined from performing FFTs for the data in panels a-c, with symbols plotted in the same colors as the waveforms. Only fundamental components are included. Good correspondence is observed between the three measurements. Multiple frequency branches are resolved, indicating a complex multi-sheet Fermiology.
Figure 2: Temperature evolution of (left) quantum oscillatory waveforms, (centre) their corresponding FFT spectra and (right) Lifshitz-Kosevich fits to the $T$-dependence of FFT amplitudes for (a) torque measurements with H $\parallel$ [101], (b) torque measurements with H tilted three degrees away from [101]* towards [100] and (c) TDO measurements with H $\parallel$ [101]*. Multiple frequency branches are observed ranging from 900 T to 8.3 kT, with a spread of effective quasiparticle masses up to 4.8(3) $m_{e}$.
@@ III Results
Fig. 1 presents QOs from our SECUF, HFML-EMFL and WNHMFC experiments, for rotations of the orientation of H through the [001]-[101]-[100] crystallographic plane. Fig. 1e collates the results of these experiments by plotting the oscillatory frequencies versus rotation angle, where the frequencies have been computed by fast Fourier transforms (FFTs) performed on the QO traces at each angle. Good correspondence between the three datasets is observed, with several frequency branches clearly resolved.
In Appendix A we examine the harmonic content of the quantum oscillatory frequency spectra. The pristine quality of single crystal specimens measured in this study is underlined by the observation of an eighth harmonic component at low temperatures and high magnetic fields, at a frequency of 71.5 kT, indicative of a mean free path $\gtrsim$ 0.5 $\upmu$m. This tallies with the residual resistivity of 0.1 $\upmu\Omega$cm, which by simple Drude considerations implies a mean free pah of $\approx 0.9$ $\upmu$m. Throughout the present section we focus solely on the fundamental QO frequency components.
In Fig. 2 we plot the evolution in temperature $T$ of the oscillatory waveforms for torque measured with H aligned along the [101] direction and close to the [101]* direction, and for TDO with H $\parallel$ [101]*. (See Appendix B for a discussion of the crystal orientation.) Numerous frequency components with a range of effective cyclotron masses are discerned. Good correspondence is observed between the frequency spectra and corresponding effective masses resolved in both our magnetic torque and contactless resistivity experiments. As $\bm{\tau}$ is a bulk thermodynamic property of the material, this close correspondence indicates that our contactless resistivity measurements are sensitive to the bulk band structure of RuO2.
@@
Prior density functional theory (DFT) calculations [35] predicted that the observed QO frequencies in the [100]-[010] rotation plane should yield starkly different angular profiles depending on whether RuO2 is antiferromagnetic or paramagnetic. Fig. 3 compares the angular dependence of QO frequencies observed by experiment with the DFT predictions for nonmagnetic and AM ordering reported in ref. [35], along with the results of our own DFT calculations (see Appendix C for details). While the predicted AM Fermi surface from ref. [35] would not be expected to yield any frequency components $>3$ kT in this rotation plane, the nonmagnetic Fermi surfaces from that study and from our calculations should both manifest a range of QO frequencies including a prominent branch at $\approx 9$ kT. This prediction matches well with what we observe experimentally. In Fig. 3c we resolve numerous frequency components ranging from 850 T to 11 kT.
@@
Figure 4: Angular evolution of quantum oscillatory frequencies in the [001]-[100] rotation plane. (a) Fundamental QO frequencies versus rotation angle $\theta$ reproduced from Fig. 1. Here $\theta=0\degree$ corresponds to H $\parallel$ [001] and $\theta=90\degree$ to H $\parallel$ [100]. (b) Simulated QO frequencies versus $\theta$ computed from our DFT calculations for a nonmagnetic Fermi surface. (c) The predicted angular frequency distribution in the absence of magnetism from ref. [35] and (d) the alternative prediction for the case of AM ordering. (e) Rendering of the nonmagnetic RuO2 Fermi surface from our DFT calculations. The red, blue and green coloring throughout this figure is indicative of the three separate Fermi sheets.
@@ IV Discussion
If AM ordering were to be manifested throughout the bulk of RuO2, a markedly different Fermi surface geometry and topology has been expected [2, 35] compared to the nonmagnetic case. The DFT calculations of both the present study and ref. [35], for the scenario of no magnetic ordering in RuO2, capture the main qualitative features of the quantum oscillatory frequency spectra resolved by experiment. Our data are well described by three Fermi surface sheets (Fig. 4). These include two largely isotropic, almost spherical pockets, one yielding QO frequencies $\approx$ 8-9 kT, and another much smaller one at $\approx$ 1 kT. The other Fermi surface sheet (colored blue in Fig. 4e) possesses a much more complex geometry, consisting of interconnected body and neck sections with numerous extremal orbital areas contributing to the observed QO frequency spectra. We note that this interpretation of three quite distinct Fermi sheets is broadly consistent with much earlier work on RuO2 [37, 38, 39, 40, 41], in which this material was also considered to be nonmagnetic.
@@
We note further that recent photoemission experiments have discerned the presence of electronic states that appear to be confined to the surface of RuO2 with low dispersion characteristic of low Fermi velocities [24, 25]. However, we do not resolve any signature of these states in either our dHvA or SdH measurements. That torque magnetometry is not sensitive to these surface states is unsurprising, since this technique probes a bulk thermodynamic property. Consider for example a cubic sample geometry – the ratio of surface to bulk unit cells in such a system is 1:106, therefore the contributory oscillatory amplitude from any surface component would be expected to be orders of magnitude smaller than that from the bulk [42]. By contrast, contactless resistivity measurements by the TDO and PDO techniques are much more sensitive to surface effects, as they operate at MHz frequencies for which the skin depth of a high conductivity metal such as RuO2 is very short. However, the clearest photoemission spectra of the surface states appear to resolve a largely circular pocket that encloses an area (in reciprocal space) of $\approx 1$-2 [24, 25], which for H $\parallel$ [110] would yield QOs with a frequency $\approx$ 10 kT. At $\mu_{0}H=$ 30 T, the corresponding cyclotron orbit of such a pocket would possess a diameter (in real space) of $\approx$ 200 nm, necessitating a long mean free path in the surface layer. While the bulk of our samples indeed possesses such a long mean free path (see Appendix A), if the surface states are topological as has been proposed [25] then they will be confined to a very small volumetric fraction of the sample right at the surface. It would be reasonable to assume that the mean free path at the surface is less than in the bulk, which would in turn explain why we did not appear to resolve any signature of these surface states. Additionally, we note that some of the surface sheets appear to be quasi-1D, which in a magnetic field would yield open orbits that do not produce QOs.
@@
The pronounced disagreement between the expected geometry of the AM Fermi surface of RuO2 and our experimental observations gives strong corroboration to the growing body of evidence in favor of RuO2 not being magnetically ordered [21, 22, 24, 25, 43, 44]. Reconciling a nonmagnetic ground state with observations of phenomena including the anomalous Hall effect, spin-current generation and magnetic circular dichroism [15, 16, 17, 18, 19, 20] therefore presents an outstanding challenge. The anomalous character of the electronic surface states resolved by photoemission spectroscopy – reported to display a Rashba-like spin-splitting [24] with non-trivial topology [25] – may be the cause of several of these observations. It has also been proposed that RuO2 thin films may exhibit a different magnetic character compared to single crystals [45], and that the bulk magnetic properties may be acutely sensitive to the density of ruthenium vacancies [46]. We note that the observation of very high frequency QO components, with multiple harmonics up to 71.5 kT, implies a long mean free path and thus a very low density of vacancies in our measured crystals.
@@
In conclusion, we performed a quantum oscillation study of the metallic altermagnet candidate RuO2. We resolved a Fermi surface geometry that matches well with the predicted electronic structure in the absence of magnetism - by contrast, the expected angular dependence of quantum oscillatory frequency spectra for the scenario of altermagnetic ordering disagrees markedly with our experimental observations. Our findings strongly support the scenario that the magnetic ground state of RuO2 is paramagnetic in character.
@@ Acknowledgements.
We gratefully acknowledge stimulating discussions with A. Agarwal, D. Calugaru, D. Chichinadze, B. Ramshaw, T. Sato D. Shaffer, R.-J. Slager and S. Souma. This project was supported by the EPSRC of the UK (grant no. EP/X011992/1). We acknowledge the support of HFML-RU/NWO-I, member of the European Magnetic Field Laboratory (EMFL) and the Engineering and Physical Sciences Research Council (EPSRC, UK) via its membership to the EMFL (grant EP/N01085X/1). A portion of this work was carried out at the Synergetic Extreme Condition User Facility (SECUF, https://cstr.cn/31123.02.SECUF). This work was supported by JSPS KAKENHI grant no. JP22H01168. A portion of this work was performed at the National High Magnetic Field Laboratory, which is supported by National Science Foundation Cooperative Agreement No. DMR-2128556 and the State of Florida. S.P. acknowledges support of the JST Sakura Science Program. T.I.W. and A.G.E. acknowledge support from QuantEmX grants from ICAM and the Gordon and Betty Moore Foundation through Grant GBMF9616 and from the US National Science Foundation (NSF) Grant Number 2201516 under the Accelnet program of Office of International Science and Engineering (OISE). A.G.E. acknowledges support from Sidney Sussex College (University of Cambridge).

### v2

@@ Abstract
The metallic oxide RuO2 has emerged as a promising altermagnet candidate, owing to reports of this material hosting antiferromagnetic ordering accompanied by a spin-split electronic band structure characteristic of time-reversal symmetry-breaking. However, recent studies have robustly questioned this scenario. Here we map the Fermi surface of pristine single-crystalline RuO2. By measuring magnetic quantum oscillations of a bulk thermodynamic property, our study resolves the electronic structure present in the bulk of RuO2. Several Fermi sheets are discerned, with a range of effective quasiparticle masses up to five times that of the bare electron mass. We compare our measurements with the predictions for altermagnetic and nonmagnetic Fermi surfaces deduced from density functional theory calculations. The quantum oscillatory frequency spectra correspond very poorly to the profile expected for the case of altermagnetism; by contrast, they correspond well to the nonmagnetic scenario. Our findings place significant constraints on the bulk magnetic properties of RuO2, and strongly suggest that this material is a paramagnet.
@@ I Introduction
The recently proposed phenomenon of an altermagnetic (AM) phase of matter – characterized by a compensated collinear magnetic structure with a spin polarization that alternates through both the crystal structure in real space and the electronic band structure in reciprocal space – has attracted widespread attention [1, 2, 3, 4, 5, 6, 7, 8, 9, 10]. In addition to being of great fundamental interest – constituting a novel, distinct class of magnetic order – there are numerous properties of AM materials that make them highly desirable for technological applications [11, 12]. For example, their time-reversal symmetry-broken electronic band structures may, as per ferromagnetic materials, manifest spin-polarized currents accompanied by the anomalous Hall effect [4, 3]. Furthermore, attempts to miniaturize ferromagnetic components within integrated circuits often suffer deleterious effects from stray fields causing interference between neighboring elements. By contrast, the lack of a net magnetization M for AM components could enable them to be spatially positioned much closer together in next-generation spintronic circuitry.
@@ II Experimental techniques
For torque magnetometry measurements, samples were aligned by Laue diffractometry and affixed to a BeCu cantilever using a low-temperature adhesive varnish. This technique constitutes a bulk-sensitive thermodynamic measurement that provides a highly sensitive probe of the anisotropy in a material’s magnetic susceptibility tensor. Due to its high fidelity, torque magnetometry is often deployed for dHvA effect measurements in high magnetic fields [30, 31, 32, 33]. In HFML-EMFL the magnetic torque $\bm{\tau}=\mu_{0}(\textbf{M}\crossproduct\textbf{H})$ was measured capacitively by a General Radio analogue capacitance bridge, and calibrated to absolute units using an Andeen-Hagerling (AH) digital capacitance bridge. At SECUF, we measured $\tau$ directly with an AH bridge. In both sets of torque experiments, samples were rotated in-situ, with the orientation calibrated by a Hall sensor. For contactless resistivity measurements, samples were also aligned by Laue diffractometry, and affixed to hand-wound copper coils by low-temperature adhesive varnish. TDO and PDO measurements were obtained utilizing a similar methodology to that outlined in ref. [34]. At WHMFC, samples were also rotated in-situ, with the orientation determined by a pick-up coil.
Figure 1: Angle-dependent quantum oscillation measurements in RuO2. Background-subtracted magnetic torque ($\Delta\tau$) measured at incremental angles in the [001]—[100] rotation plane recorded at (a) SECUF and (b) HFML-EMFL. (c) Quantum oscillations in the contactless resistivity measured at WNHMFC, also in the [001]—[100] plane. (d) Fast Fourier transforms (FFTs) of the data from panel a. These spectra have been renormalized and offset for ease of presentation. (e) Angular distribution of quantum oscillatory frequency spectra, determined from performing FFTs for the data in panels a-c, with symbols plotted in the same colors as the waveforms. Only fundamental components are included. Good correspondence is observed between the three measurements. Multiple frequency branches are resolved, indicating a complex multi-sheet Fermiology.
Figure 2: Temperature evolution of (left) quantum oscillatory waveforms in RuO2, (centre) their corresponding FFT spectra and (right) Lifshitz-Kosevich fits to the $T$-dependence of FFT amplitudes for (a) torque measurements with H $\parallel$ [101], (b) torque measurements with H tilted three degrees away from [101]* towards [100] and (c) TDO measurements with H $\parallel$ [101]*. Multiple frequency branches are observed ranging from 900 T to 8.3 kT, with a spread of effective quasiparticle masses up to 4.8(3) $m_{e}$.
@@ III Results
Fig. 1 presents QOs from our SECUF, HFML-EMFL and WNHMFC experiments, for rotations of the orientation of H through the [001]-[100] rotation plane. Fig. 1e collates the results of these experiments by plotting the oscillatory frequencies versus rotation angle, where the frequencies have been computed by fast Fourier transforms (FFTs) performed on the QO traces at each angle. Good correspondence between the three datasets is observed, with several frequency branches clearly resolved.
In Appendix A we examine the harmonic content of the quantum oscillatory spectra. The pristine quality of single crystal specimens measured in this study is underlined by the observation of an eighth harmonic component at low temperatures and high magnetic fields, at a frequency of 71.5 kT, indicating a mean free path $\gtrsim$ 0.5 $\upmu$m. This tallies with the residual resistivity of 0.1 $\upmu\Omega$cm, which by simple Drude considerations implies a mean free path of $\approx 0.9$ $\upmu$m. Throughout the present section we focus solely on the fundamental QO frequency components.
In Fig. 2 we plot the evolution in temperature $T$ of the oscillatory waveforms for torque measured with H aligned along the [101] direction and close to the [101]* direction, and for TDO with H $\parallel$ [101]*. (See Appendix B for a discussion of the crystal orientation.) Numerous frequency components with a range of effective cyclotron masses are resolved. Good correspondence is observed between the frequency spectra and effective masses resolved in both our magnetic torque and contactless resistivity experiments. As $\bm{\tau}$ is a bulk thermodynamic property of the material, this close correspondence indicates that our contactless resistivity measurements are sensitive to the bulk band structure of RuO2.
@@
Prior density functional theory (DFT) calculations [35] predicted that the observed QO frequencies should yield starkly different angular profiles depending on whether RuO2 is altermagnetic or paramagnetic. Fig. 3 compares the angular dependence of QO frequencies observed by experiment with the DFT predictions for nonmagnetic and AM ordering reported in ref. [35], along with the results of our own DFT calculations (see Appendix C for details). While the predicted AM Fermi surface from ref. [35] would not be expected to yield any frequency components $>3$ kT in this rotation plane, the nonmagnetic Fermi surfaces from that study and from our calculations should both manifest a range of QO frequencies including a prominent branch at $\approx 9$ kT. This prediction matches well with what we observe experimentally. In Fig. 3c we resolve numerous frequency components ranging from 850 T to 11 kT.
@@
Figure 4: Angular evolution of quantum oscillatory frequencies of RuO2 in the [001]-[100] rotation plane. (a) Fundamental QO frequencies versus rotation angle $\theta$ reproduced from Fig. 1. Here $\theta=0\degree$ corresponds to H $\parallel$ [001] and $\theta=90\degree$ to H $\parallel$ [100]. (b) Simulated QO frequencies versus $\theta$ computed from our DFT calculations for a nonmagnetic Fermi surface. (c) The predicted angular frequency distribution in the absence of magnetism from ref. [35] and (d) the alternative prediction for the case of AM ordering. (e) Rendering of the nonmagnetic RuO2 Fermi surface from our DFT calculations. The red, blue and green coloring throughout this figure is indicative of the three separate Fermi sheets.
@@ IV Discussion
If AM ordering were to be manifested throughout the bulk of RuO2, a markedly different Fermi surface geometry and topology has been expected [2, 35] compared to the nonmagnetic case. The DFT calculations of both the present study and ref. [35] for the scenario of no magnetic ordering in RuO2 capture the main features of the quantum oscillatory frequency spectra resolved by experiment. Our data are well described by three Fermi surface sheets (Fig. 4). These include two largely isotropic, almost spherical pockets, one yielding QO frequencies $\approx$ 8-9 kT, and another much smaller one at $\approx$ 1 kT. The other Fermi surface sheet (colored blue in Fig. 4e) possesses a much more complex geometry, consisting of interconnected body and neck sections with numerous extremal orbital areas contributing to the observed QO frequency spectra. We note that this interpretation of three quite distinct Fermi sheets is broadly consistent with much earlier work on RuO2 [37, 38, 39, 40, 41], in which this material was also considered to be nonmagnetic.
@@
Recent photoemission experiments have discerned the presence of electronic states that appear to be confined to the surface of RuO2 with low dispersion characteristic of low Fermi velocities [24, 25]. However, we do not resolve any signature of these states in either our dHvA or SdH measurements. That torque magnetometry is not sensitive to these surface states is unsurprising, since this technique probes a bulk thermodynamic property. Consider for example a cubic sample geometry – the ratio of surface to bulk unit cells in such a system is 1:106, therefore the contributory oscillatory amplitude from any surface component may be expected to be orders of magnitude smaller than that from the bulk [42]. By contrast, contactless resistivity measurements by the TDO and PDO techniques are much more sensitive to surface effects, as they operate at MHz frequencies for which the skin depth of a high conductivity metal such as RuO2 is very short. However, the clearest photoemission spectra of the surface states appear to resolve a largely circular pocket that encloses an area (in reciprocal space) of $\approx 1$-2 [24, 25], which for H $\parallel$ [110] would yield QOs with a frequency $\approx$ 10 kT. At $\mu_{0}H=$ 30 T, the corresponding cyclotron orbit of such a pocket would possess a diameter (in real space) of $\approx$ 200 nm, necessitating a long mean free path in the surface layer. While the bulk of our samples indeed possesses such a long mean free path (see Appendix A), if the surface states are topological as has been proposed [25] then they will be confined to a very small volumetric fraction of the sample right at the surface. It would be reasonable to assume that the mean free path at the surface is less than in the bulk, which would in turn explain why we did not appear to resolve any signature of these surface states. Additionally, we note that some of the surface sheets appear to be quasi-1D, which in a magnetic field would yield open orbits that do not produce QOs.
@@
The pronounced disagreement between the expected geometry of the AM Fermi surface of RuO2 and our experimental observations corroborates the growing body of evidence in favor of RuO2 not being magnetically ordered [21, 22, 24, 25, 43, 44]. Reconciling a nonmagnetic ground state with phenomena including the anomalous Hall effect, spin-current generation and magnetic circular dichroism [15, 16, 17, 18, 19, 20] therefore presents an outstanding challenge. The anomalous character of the electronic surface states resolved by photoemission spectroscopy – reported to display a Rashba-like spin-splitting [24] with non-trivial topology [25] – may be the cause of several of these observations. It has also been proposed that RuO2 thin films may exhibit a different magnetic character compared to single crystals [45], and that the bulk magnetic properties may be acutely sensitive to the density of ruthenium vacancies [46]. We note that the observation of very high frequency QO components, with multiple harmonics up to 71.5 kT, implies a long mean free path and thus a very low density of vacancies in our measured crystals.
@@
In conclusion, we performed a quantum oscillation study of the metallic altermagnet candidate RuO2. We resolved a Fermi surface geometry that matches well with the predicted electronic structure in the absence of magnetism; whereas, the expected angular dependence of quantum oscillatory frequency spectra for the case of altermagnetic ordering disagrees markedly with our experimental observations. Our findings strongly support the scenario that the magnetic ground state of RuO2 is paramagnetic in character.
@@ Acknowledgements.
We gratefully acknowledge stimulating discussions with A. Agarwal, D. Calugaru, D. Chichinadze, A. Coldea, B. Ramshaw, T. Sato, D. Shaffer, R.-J. Slager and S. Souma. This project was supported by the EPSRC of the UK (grant no. EP/X011992/1). We acknowledge the support of HFML-RU/NWO-I, member of the European Magnetic Field Laboratory (EMFL) and the Engineering and Physical Sciences Research Council (EPSRC, UK) via its membership to the EMFL (grant EP/N01085X/1). A portion of this work was carried out at the Synergetic Extreme Condition User Facility (SECUF, https://cstr.cn/31123.02.SECUF). This work was supported by JSPS KAKENHI grant no. JP22H01168. A portion of this work was performed at the National High Magnetic Field Laboratory, which is supported by National Science Foundation Cooperative Agreement No. DMR-2128556 and the State of Florida. S.P. acknowledges support of the JST Sakura Science Program. T.I.W. and A.G.E. acknowledge support from QuantEmX grants from ICAM and the Gordon and Betty Moore Foundation through Grant GBMF9616 and from the US National Science Foundation (NSF) Grant Number 2201516 under the Accelnet program of Office of International Science and Engineering (OISE). A.G.E. acknowledges support from Sidney Sussex College (University of Cambridge).